7^(x+8)=6^x
(the answer must be rounded to three decimal places).

Answer by stanbon(75887)

7^(x+8)=6^x
---
Take the natural log of both sides to get:
(x+8)*ln(7) = x*ln(6)
---
(ln(7)-ln(6))x = -8ln(7)
---
0.15415*x = -15.5673
x = -100.99
=====================
Cheers,
Stan H.
=====================